# Strong Haken via Thin Position

### v1

@@ Abstract.
We use thin position of Heegaard splittings to give a new proof of Haken’s Lemma that a Heegaard surface of a reducible manifold is reducible and of Scharlemann’s “Strong Haken Theorem”, that a Heegaard surface for a 3-manifold may be isotoped to intersect a given collection of essential spheres and discs in a single loop each. This article could also serve as an introduction to the theory of generalized Heegaard surfaces and it includes a careful study of their behaviour under amalgamation.
@@ 1. Introduction
A surface $H$ in a closed, connected, orientable 3-manifold $M$ is a Heegaard surface if $M$ can be obtained (up to homeomorphism) by gluing together two handlebodies via a homeomorphism of their boundaries such that $H$ is the image of the boundaries. This definition can be generalized to 3-manifolds with boundary, by allowing $H$ to divide $M$ into two compressionbodies, rather than handlebodies. The Heegaard genus $\operatorname{g}(M)$ of $M$ is the minimal genus of a Heegaard surface in $M$. A 3-manifold $M$ is reducible if there is an embedded sphere $S\subset M$ such that $S$ is not the boundary of a 3-ball in $M$. A reducing sphere for a Heegaard surface is an essential sphere $S\subset M$ such that $S\cap H$ consists of a single curve in $H$. A Heegaard surface is reducible if there exists a reducing sphere.
@@
Lemma 1.1 (Haken’s Lemma (1968) [Haken]).
A Heegaard surface $H$ for a compact, orientable reducible 3-manifold $M$ is reducible.
@@
Corollary 1.2 (Heegaard genus is additive).
@@
Casson and Gordon [CG] adapted Haken’s result so that it also applies to discs:
@@
If $H$ is a Heegaard surface for a compact orientable 3-manifold $M$ with boundary such that $M$ contains an essential disc $D$, then there exists a disc $D^{\prime}$ with the same boundary as $D$, such that $D$ and $H$ intersect in a single simple closed curve.
@@
Noting that Haken’s result only guarantees existence, in 2020, Scharlemann considered the question of whether if $S$ is a fixed sphere or disc, a Heegaard surface $H$ can necessarily be isotoped to intersect the given sphere or disc $S$ in a single simple closed curve. He showed:
Theorem 1.5 (Strong Haken Theorem [Scharlemann]).
Suppose that $\mathcal{S}$ is the union of pairwise disjoint essential spheres and discs in a compact, connected, orientable 3-manifold $M$. If $H$ is a Heegaard surface for $M$, then $H$ can be isotoped so that $H$ intersects each component of $\mathcal{S}$ in a single simple closed curve.
@@ 2. Outline
Heegaard splitting splitting theory has reached considerable maturity as a mathematical subject. A key insight that is the basis of many modern results is the realization (perhaps originating in [G3, Kobayashi]) that strongly irreducible Heegaard splittings can isotoped to intersect essential surfaces in circles of intersection that are mutually essential or mutually inessential. This result does not use Haken’s Lemma; we sketch the proof in Lemma 3.6. Before expounding on some of the details, here is a brief sketch of the argument, for the experts,. For simplicity, assume that $H$ is a Heegaard surface for a closed, connected, orientable, reducible 3-manifold $M=M_{1}\#M_{2}$ where both $M_{1}$ and $M_{2}$ are prime and irreducible. Let $S\subset M$ be an essential sphere. In this simplified setting $S$ is unique up to isotopy. (Consequently, Haken’s Lemma will directly imply the Strong Haken Theorem, however to give an argument that applies more generally we won’t use that fact.) Thin $H$ to a generalized Heegaard surface $\mathcal{H}$ consisting of thick surfaces $\mathcal{H}^{+}$ and thin surfaces $\mathcal{H}^{-}$. This can be done so that $\mathcal{H}^{-}$ is incompressible and each component of $\mathcal{H}^{+}$ is strongly irreducible. The incompressibility of $\mathcal{H}^{-}$ and the strong irreducibility of $\mathcal{H}^{+}$ can be used to isotope $\mathcal{H}$ to be disjoint from $S$. It then follows that some component of $\mathcal{H}^{-}$ is an essential sphere $P$. By our simplifications, the surface $S$ must be parallel to $P$ and surgery on $P$ produces $M_{1}$ and $M_{2}$. On each side of $P$ amalgamate $\mathcal{H}$ to Heegaard surfaces $H_{1}$ and $H_{2}$ for $M_{1}$ and $M_{2}$. We may then amalgamate $H_{1}\cup P\cup H_{2}$ to a Heegaard surface $H^{\prime}$. We perform this amalgamation carefully. Start by sliding the feet of the 1-handles, of the compressionbody between $P$ and $H_{1}$ so that when we extend them to lie on $H_{2}$, we need only extend a single foot. This ensures that $P$ (and, in this simplified setting, $S$) intersect $H^{\prime}$ in a single simple closed curve. We may then appeal to a result of Lackenby [Lackenby, Prop. 3.1] which guarantees that $H$ is isotopic to $H^{\prime}$. In the general setting, where $M$ admits multiple essential spheres and discs, and especially when some of those spheres may be nonseparating, we need to do more work to implement the concept.
@@
Here is a bit more information on the overall strategy. A Heegaard surface is weakly reducible if it admits disjoint compressing discs on opposites sides. A Heegaard surface is strongly irreducible if it is not weakly reducible. Thin position for 3-manifolds is technique for acquiring topological information from a weakly reducible Heegaard splitting of a 3-manifold $M$. It involves replacing a Heegaard surface with a (likely disconnected surface) called a generalized Heegaard surface; each component of a generalized Heegaard surface is either a thick surface or a thin surface. The thick surfaces are Heegaard surfaces for the regions of $M$ complementary to the thin surfaces. The basic idea (originating with Casson and Gordon [CG]) is to decompose a weakly reducible Heegaard surface into three mutually disjoint surfaces (each is potentially disconnected), two of which are thick and one of which is thin. Thinking of a Heegaard splitting as a sequence of handle attachments, where 0-handles and 1-handles are attached before 2-handles and 3-handles, this operation (called “weak reduction”) is equivalent to attaching a 2-handle (and possibly 3-handle) prior to attaching a 1-handle (and possibly a 0-handle). If one or both of the thick surfaces are themselves weakly reducible we can repeat the operation. The end result is a collection of thick and thin surfaces. If it is ever the case that a thick and thin surface are parallel, we remove them from the collection. This way of creating and removing thick and thin surfaces is called thinning. Scharlemann and Thompson [ST] produce a complexity that decreases under thinning; a collection of thin and thick surfaces is thin if it minimizes their complexity. In a thin collection, the thin surfaces are incompressible and the thick surfaces are strongly irreducible. (In [ST, Rule 5], the authors appeal to Lemma 1.3 for this fact; we replace it with Lemma 3.6.) Scharlemann and Thompson observe that their complexity is (in a certain sense) additive under connected sum. Although they do not indicate what they had in mind for a proof, it is likely they intended to use Haken’s Lemma, together with the prime decomposition theorem for 3-manifolds. Our route on the other hand, proves Haken’s Lemma and the Strong Haken Theorem using (in a certain sense) additivity.
@@
The proof of the Strong Haken Theorem is then carried out at the end of Section 6 and in Section 7 as follows. We are given a Heegaard surface $H$ for a 3-manifold that contains essential spheres and discs. We are also given a mutually disjoint collection $\mathcal{S}$ of essential spheres and discs. We thin $H$ to arrive at a thin generalized Heegaard surface $\mathcal{H}$. We use juggling to ensure it is disjoint from $\mathcal{S}$. We then carefully amalgamate $\mathcal{H}$ to a Heegaard surface $J$ and appeal to Lackenby’s result to know that $H$ is isotopic to $J$.
@@ 2.1. A word on notation
Unless otherwise specified all 3-manifolds $M$ are compact, connected, and orientable. All surfaces are compact and orientable. In a directed graph (aka digraph) a coherent cycle (resp. coherent path) is an embedded oriented loop (resp. path) such that the orientation of the loop (resp. path) coincides with the orientation of each edge it traverses. As in graph theory, a digraph is acyclic if it does not contain a coherent cycle; the underlying undirected graph of an acyclic digraph may contain cycles. We let $I=[0,1]$. If $D$ is the union of pairwise disjoint discs properly embedded in $M$, we let $M|_{D}$ denote the result of $\partial$-reducing $M$ along $D$. The scars from $D$ are the discs in $\partial M|_{D}$ that are the feet of the 1-handles with cocores $D$ removed from $M$. Similarly, if $S\subset M$ is a sphere, we let $M|_{S}$ denote the result of surgery along $S$, that is the result of removing a regular neighborhood of $S$ from $M$ and then gluing in 3-balls to the resulting boundary components. In $M\setminus S$, the scars from $S$ are the feet of the copies of $S^{2}\times I$ removed from $M$. In $M|_{S}$, the scars from $S$ are the 3-balls that were glued to the scars in $M\setminus S$.
@@
An s-compressing disc is a generalization of “compressing disc”: the definition is identical except we replace (3) with:
@@
A semi-compressing disc for $S$ is an s-compressing disc that is not a compressing disc. It necessarily has inessential boundary in $S$. If $\Delta$ is the union of mutually disjoint s-discs for $S$, we can compress $S$ to a (possibly disconnected) surface $S|_{\Delta}$ by letting $S|_{\Delta}$ be the result of removing a regular neighborhood $N(\Delta)\approx(\Delta\times I)$ of $\partial\Delta$ from $S$ and then taking the union with $\Delta\times\partial I$. Observe that $S|_{\Delta}$ inherits a normal orientation from $S$ and that each disc in $\Delta$ leaves two scars on $S|_{\Delta}$.
@@
  • compressible if there is a compressing disc for $S$;
  • weakly reducible if there exist disjoint s-discs $D_{1}$ and $D_{2}$ for $S$, with boundaries on the same component of $S$, such that near $S$, the discs $D_{1}$ and $D_{2}$ lie on opposite sides of the component of $S$ containing their boundaries and at least one of $D_{1}$ or $D_{2}$ is a compressing disc.
@@
    There is a (possibly empty) collection $\Delta$ of mutually disjoint s-discs for $\partial_{+}C$, such that $C|_{\Delta}$ is the union of trivial compressionbodies and with the designations of $\partial_{+}$ and $\partial_{-}$ also inherited from $C$. (Such a collection $\Delta$ is a complete collection of s-discs for $C$.
A punctured trivial compressionbody is the result of removing the open neighborhood of a finite (possibly empty) collection of points from the interior of a trivial compressionbody. The positive boundary of a punctured product compressionbody is the same as the positive boundary prior to removing the open balls. Observe that a punctured product compressionbody is a compressionbody. If a compressionbody does not admit any compressing discs for its positive boundary it is a punctured trivial compressionbody.
@@
    $C$ does not contain an s-disc for $\partial_{+}C$ if and only if it is a ball compressionbody or product compressionbody. If $C$ contains an s-disc for $\partial_{+}C$ but does not contain a compressing disc for $\partial_{+}C$, then it is a punctured product compressionbody.
@@
By definition of semi-compressing disc, $\partial D$ bounds at least one disc in $\partial_{-}C$. As every sphere in a compressionbody is separating and since $\partial_{-}C$ is incompressible in $C$, there is such a disc $E$ such that the sphere $D\cup E$ bounds a submanifold $W\subset C$ disjoint from $\partial_{-}C$. An innermost disc argument ensures that there is a maximal pairwise disjoint collection of compressing discs for $\partial_{+}C$ disjoint from $D$. Thus, after capping off spherical components of $\partial_{-}C$ (other than the one containing $\partial D$ if it is a sphere) we see that $W$ becomes a 3-ball. Thus, reconstructing $C$, we see that $D\cup E$ bounds a compressionbody in $C$ that is the result of removing at least one open 3-ball from a 3-ball. ∎
@@
Given any compressionbody $C$ with $\partial_{-}C\neq\varnothing$, there exists a collection of s-discs $D\subset C$ such that each component of $C|_{D}$ is either a handlebody or a product compressionbody and each component contains exactly one scar from $D$. Furthermore, if $A\subset C$ is the union of pairwise disjoint vertical annuli and compressing discs for $\partial_{+}C$, we may choose $D$ to be disjoint from $A$.
@@
Figure 2. By sliding 1-handles, we may find discs separating a compressionbody into product compressionbodies and handlebodies so that each product compressionbody contains a single scar.
@@
where the unions appearing in the last two items are over all the components of $M\setminus\mathcal{H}$ and the equality holds after gluing. A generalized Heegaard surface $\mathcal{H}$ is oriented if each component of $\mathcal{H}\cup\partial M$ is assigned a transverse orientation such that for each compressionbody $C\subset M\setminus\mathcal{H}$, $C$ is a cobordism from $\partial_{-}C$ to $\partial_{+}C$ or vice-versa. We say that a compressionbody $C$ is above (resp. below) $\partial_{+}C$ if the transverse orientation for $\partial_{+}C$ points into (resp. out of) $C$. Two oriented generalized Heegaard surfaces are equivalent if they are ambient isotopic in $M$, via an isotopy preserving orientations. We will only work with oriented generalized Heegaard surfaces in what follows.
@@
A generalized Heegaard surface is a Heegaard surface if it is connected. In which case, $\mathcal{H}=\mathcal{H}^{+}$ and $\mathcal{H}^{-}=\varnothing$.

### v2

@@ Abstract.
We use thin position of Heegaard splittings to give a new proof of Haken’s Lemma that a Heegaard surface of a reducible manifold is reducible and of Scharlemann’s “Strong Haken Theorem”: a Heegaard surface for a 3-manifold may be isotoped to intersect a given collection of essential spheres and discs in a single loop each. We also give a reformulation of Casson and Gordon’s theorem on weakly reducible Heegaard splittings, showing that they exhibit additional structure with respect to certain incompressible surfaces. This article could also serve as an introduction to the theory of generalized Heegaard surfaces and it includes a careful study of their behaviour under amalgamation.
@@ 1. Introduction
A surface $H$ in a closed, connected, orientable 3-manifold $M$ is a Heegaard surface if $M=H_{1}\cup_{H}H_{2}$ where $H_{1}$ and $H_{2}$ are handlebodies and (after gluing) $H=\partial H_{1}=\partial H_{2}$. This definition can be generalized to 3-manifolds with boundary, by allowing $H$ to divide $M$ into two compressionbodies, rather than handlebodies. The Heegaard genus $\operatorname{g}(M)$ of $M$ is the minimal genus of a Heegaard surface in $M$. A 3-manifold $M$ is reducible if there is an embedded sphere $S\subset M$ such that $S$ is not the boundary of a 3-ball in $M$; the sphere $S$ is a reducing sphere. A Heegaard surface is reducible if there exists a reducing sphere intersecting it in a single simple closed curve.
@@
Theorem (Haken’s Lemma (1968) [Haken]).
Every Heegaard surface $H$ for a compact, connected, orientable reducible 3-manifold $M$ is reducible.
@@
Corollary (Heegaard genus is additive).
@@
A sphere $S\subset M$ is essential if it is a reducing sphere that is not $\partial$-parallel in $M$. If $\partial M$ does not contain a sphere, then every sphere in $M$ is a reducing sphere if and only if it is essential. If $M$ is not a 3-ball but $\partial M$ contains a sphere, then any $\partial$-parallel sphere is also a reducing sphere. To avoid trivialities, we will generally phrase our results for essential spheres, rather than reducing spheres. A properly embedded disc $D$ in $M$ is essential if $\partial D$ does not bound a disc in $\partial M$. Casson and Gordon [CG] adapted Haken’s result so that it also applies to discs:
@@
If $H$ is a Heegaard surface for a compact orientable 3-manifold $M$ with boundary such that $M$ contains an essential disc $D$, then there exists an essential disc $D^{\prime}$ with the same boundary as $D$, such that $D$ and $H$ intersect in a single simple closed curve.
@@
Noting that Haken’s result only guarantees existence, in 2020, Scharlemann considered the question of whether if $S$ is a fixed essential sphere or disc, a Heegaard surface $H$ can necessarily be isotoped to intersect the given sphere or disc $S$ in a single simple closed curve. He showed:
Theorem (Strong Haken Theorem [Scharlemann]).
Suppose that $\mathcal{S}$ is the union of pairwise disjoint properly embedded essential spheres and discs in a compact, connected, orientable 3-manifold $M$. If $H$ is a Heegaard surface for $M$, then $H$ can be isotoped so that $H$ intersects each component of $\mathcal{S}$ in a single simple closed curve.
@@
Our approach is rooted in the Casson-Gordon Theorem for Heegaard surfaces [CG], and, in turn, gives a new perspective on it. A Heegaard surface is weakly reducible if it admits disjoint compressing discs on opposites sides. Such a pair of discs is a weak reducing pair. A Heegaard surface is strongly irreducible if it is not weakly reducible. A Heegaard surface $H$ is stabilized if there are compressing discs for $H$ on opposite sides of $H$ and with boundaries transversally intersecting in a single point. Such a pair of discs is a stabilizing pair.
Theorem (Casson-Gordon).
If $H$ is a weakly reducible Heegaard surface for a closed 3-manifold $M$, then either $H$ is stabilized or $M$ contains a closed essential surface.
It follows from their techniques that the closed essential surface has genus at most that of $H$. In their context, “essential” means that the surface is incompressible and not a 2-sphere bounding a 3-ball. They actually show that if $H$ is weakly reducible, then either $H$ is reducible or $M$ contains a closed essential surface. Waldhausen [Waldhausen-3sph] showed that every positive genus Heegaard surface of $S^{3}$ is stabilized; thus the existence of an inessential sphere intersecting $H$ in a single simple closed curve implies that $H$ is reducible.
The situation when $\partial M\neq\varnothing$ is somewhat more complicated and there are counter-examples to the theorem if we naively drop the requirement that $M$ be closed. The possibility of a Casson-Gordon theorem for 3-manifolds with boundary was first studied by Moriah [Moriah], Moriah-Sedgwick [MS], and Kobayashi [Kobayashi2]. Hayashi-Shimokawa [HS] improved on that work [^1] Hayashi-Shimokawa use the term netted instead of $\partial$-stabilized. and proved:
Theorem (Hayashi-Shimokawa).
If $H$ is a weakly reducible Heegaard surface for a compact, orientable 3-manifold (possibly with boundary) $M$, then either $H$ is stabilized, $H$ is $\partial$-stabilized, or $M$ contains a closed essential surface.
One way of defining $\partial$-stabilization is as follows. Let $M$ be a compact, orientable 3-manifold having a component $F\subset\partial M$ of positive genus. Let $J$ be a Heegaard surface for $M$ and let $H^{\prime}$ be a parallel copy of $F$ separating $F$ from $J$. Let $H$ be the result of tubing $J$ to $H^{\prime}$ via a tube that is vertical in the compressionbody between $J$ and $H^{\prime}$. We say that $H$ is obtained by a $\partial$-stabilization of $J$ along $F$. Alternatively, one can show that $H$ is $\partial$-stabilized along $F$ if there exists a compressing disc $D$ for $H$ such that $H|_{D}$ consists of two surfaces, one of which is a Heegaard surface for $M$ and the other is parallel to $F$. The disc $D$ is called a $\partial$-stabilizing disc for $H$.
Hayashi-Shimokawa’s theorem relies on the classification of Heegaard splittings of $(\text{surface})\times I$ by Scharlemann-Thompson [ST-classification]. Our work, also relying on the Waldhausen and Scharlemann-Thompson theorems, gives a somewhat different proof of Hayashi-Shimokawa’s theorem. Indeed, we prove the following which recasts the Casson-Gordon theorem and the Hayashi-Shimokawa theorem in the mold of Haken’s Lemma. In the following, as in the rest of the paper, an essential surface in a 3-manifold is a surface that is incompressible, not $\partial$-parallel, and not a 2-sphere bounding a 3-ball; every 2-sphere is incompressible, but perhaps not essential. The statement we give here is detailed, but hopefully of interest. See Figure 1 below for a depiction of the claim about Haken Sums.
Theorem 5.13 (Structure Theorem for Weakly Reducible Heegaard Surfaces).
Suppose that $H$ is an oriented weakly reducible Heegaard surface for a compact, orientable 3-manifold (possibly with boundary) $M$. Then there exists an oriented incompressible surface $F\subset M$ such that:
  1. (1)
    For each component $F^{\prime}$ of $F$, $F^{\prime}\cap H$ consists of a single simple closed curve, which is inessential in $F^{\prime}$. The curves of $F\cap H$ bound discs in $F$ all on the same side of $H$; we may perform the construction so that the discs lie on which ever side we wish.
  2. (2)
    $F\cap H$ separates $H$.
  3. (3)
    For each component $F^{\prime}$ of $F$, $\chi(F^{\prime})<-\chi(H)$.
  4. (4)
    There is a weak reducing pair of discs separated by $F$.
  5. (5)
    If some component of $F$ is a 2-sphere bounding a 3-ball, then there is a stabilizing pair of discs for $H$ disjoint from $F$
  6. (6)
    If some component of $F$ is $\partial$-parallel, then there is a $\partial$-stabilizing disc for $H$ disjoint from $F$.
  7. (7)
    There is a generalized Heegaard surface $\mathcal{H}$ for $M$ such that $\mathcal{H}^{+}$ is isotopic to $H+F$, the Haken sum of $H$ and $F$, and $\mathcal{H}^{-}=F$; furthermore, the thick surfaces $\mathcal{H}^{+}$ are strongly irreducible.
Figure 1. The top figure shows a schematic depiction of the Heegaard surface $H$ (in black) and the incompressible surface $F$. For both surfaces we indicate transverse orientations. In the second figure, in black, we show the Haken sum $H+F$ of $H$ and $F$; this is the surface obtained by smoothing the intersections using the orientations. According to Theorem 5.13, after a small isotopy $(H+F)\cup F$ is a generalized Heegaard surface.
@@ 2. Outline
Heegaard splitting theory has reached considerable maturity as a mathematical subject. A key insight that is the basis of many modern results is the realization (perhaps originating in [G3, Kobayashi]) that strongly irreducible Heegaard splittings can isotoped to intersect essential surfaces in circles of intersection that are mutually essential or mutually inessential. This result does not use Haken’s Lemma; we sketch the proof in Lemma 3.6. Before expounding on some of the details, here is a brief sketch of the argument proving the Strong Haken Theorem, for the experts. For simplicity, assume that $H$ is a Heegaard surface for a closed, connected, orientable, reducible 3-manifold $M=M_{1}\#M_{2}$ where both $M_{1}$ and $M_{2}$ are prime and irreducible (hence, not 3-balls). Let $S\subset M$ be an essential sphere. In this simplified setting $S$ is unique up to isotopy. (Consequently, Haken’s Lemma will directly imply the Strong Haken Theorem, however to give an argument that applies more generally we won’t use that fact.) Thin $H$ to a generalized Heegaard surface $\mathcal{H}$ consisting of thick surfaces $\mathcal{H}^{+}$ and thin surfaces $\mathcal{H}^{-}$. This can be done so that $\mathcal{H}^{-}$ is incompressible and each component of $\mathcal{H}^{+}$ is strongly irreducible. The incompressibility of $\mathcal{H}^{-}$ and the strong irreducibility of $\mathcal{H}^{+}$ can be used to isotope $\mathcal{H}$ to be disjoint from $S$. It then follows that some component of $\mathcal{H}^{-}$ is an essential sphere $P$. By our simplifications, the surface $S$ must be parallel to $P$ and surgery on $P$ produces $M_{1}$ and $M_{2}$. On each side of $P$ amalgamate $\mathcal{H}$ to Heegaard surfaces $H_{1}$ and $H_{2}$ for $M_{1}$ and $M_{2}$. We may then amalgamate $H_{1}\cup P\cup H_{2}$ to a Heegaard surface $J$. We perform this amalgamation carefully. Start by sliding the feet of the 1-handles, of the compressionbody between $P$ and $H_{1}$ so that when we extend them to lie on $H_{2}$, we need only extend a single foot. This ensures that $P$ (and, in this simplified setting, $S$) intersect $J$ in a single simple closed curve. We may then appeal to a result of Lackenby [Lackenby]*Prop. 3.1 which guarantees that $H$ is isotopic to $J$. In the general setting, where $M$ admits multiple essential spheres and discs, and especially when some of those spheres may be nonseparating, we need to do more work to implement the concept.
@@
Here is a bit more information on the overall strategy. Thin position for 3-manifolds is technique for acquiring topological information from a weakly reducible Heegaard splitting of a 3-manifold $M$. It involves replacing a Heegaard surface with a (likely disconnected surface) called a generalized Heegaard surface; each component of a generalized Heegaard surface is either a thick surface or a thin surface. The thick surfaces are Heegaard surfaces for the regions of $M$ complementary to the thin surfaces. The basic idea (originating with Casson and Gordon [CG]) is to decompose a weakly reducible Heegaard surface into three mutually disjoint surfaces (each is potentially disconnected), two of which are thick and one of which is thin. Thinking of a Heegaard splitting as a sequence of handle attachments, where 0-handles and 1-handles are attached before 2-handles and 3-handles, this operation (called “weak reduction”) is equivalent to attaching a 2-handle (and possibly 3-handle) prior to attaching a 1-handle (and possibly a 0-handle). If one or both of the thick surfaces are themselves weakly reducible we can repeat the operation. The end result is a collection of thick and thin surfaces. If it is ever the case that a thick and thin surface are parallel, we remove them from the collection. This way of creating and removing thick and thin surfaces is called thinning. Scharlemann and Thompson [ST] produce a complexity that decreases under thinning; a collection of thin and thick surfaces is thin if it minimizes their complexity. In a thin collection, the thin surfaces are incompressible and the thick surfaces are strongly irreducible. (In [ST, Rule 5], the authors appeal to Lemma 1.1 for this fact; we replace it with Lemma 3.6.) Scharlemann and Thompson observe that their complexity is (in a certain sense) additive under connected sum. Although they do not indicate what they had in mind for a proof, it is likely they intended to use Haken’s Lemma, together with the prime decomposition theorem for 3-manifolds. Our route on the other hand, proves Haken’s Lemma and the Strong Haken Theorem using (in a certain sense) additivity.
@@
Also, in Section 5 we prove the Structure Theorem for Weakly Reducible Heegaard Surfaces (Theorem 5.13).
@@
The proof of the Strong Haken Theorem is then carried out at the end of Section 6 and in Section 7 as follows. We are given a Heegaard surface $H$ for a 3-manifold that contains properly embedded essential spheres and discs. We are also given a mutually disjoint collection $\mathcal{S}$ of essential spheres and discs. We thin $H$ to arrive at a thin generalized Heegaard surface $\mathcal{H}$. We use juggling to ensure it is disjoint from $\mathcal{S}$. We then carefully amalgamate $\mathcal{H}$ to a Heegaard surface $J$ and appeal to a version of Lackenby’s result (Corollary 5.10) to know that $H$ is isotopic to $J$.
@@ 2.1. A word on notation
Unless otherwise specified all 3-manifolds $M$ are compact, connected, and orientable. All surfaces are compact and orientable; if they are said to lie in a 3-manifold they are usually assumed to be properly embedded. In a directed graph (aka digraph) a coherent cycle (resp. coherent path) is an embedded oriented loop (resp. path) such that the orientation of the loop (resp. path) coincides with the orientation of each edge it traverses. As in graph theory, a digraph is acyclic if it does not contain a coherent cycle; the underlying undirected graph of an acyclic digraph may contain cycles. We let $I=[0,1]$. For a (properly embedded) surface $S$ in a 3-manifold $M$, we let $M\setminus S$ denote the result of removing a regular neighborhood of $S$ from $M$. If $D$ is the union of pairwise disjoint discs properly embedded in $M$, we let $M|_{D}$ denote the result of $\partial$-reducing $M$ along $D$. The scars from $D$ are the discs in $\partial M|_{D}$ that are the feet of the 1-handles with cocores $D$ removed from $M$. Similarly, if $S\subset M$ is a sphere, we let $M|_{S}$ denote the result of surgery along $S$, that is the result of removing a regular neighborhood of $S$ from $M$ and then gluing in 3-balls to the resulting boundary components. In $M\setminus S$, the scars from $S$ are the feet of the copies of $S^{2}\times I$ removed from $M$. In $M|_{S}$, the scars from $S$ are the 3-balls that were glued to the scars in $M\setminus S$.
@@
An s-disc is a generalization of “compressing disc”: the definition is identical except we replace (3) with:
@@
A semi-compressing disc for $S$ is an s-disc that is not a compressing disc. It necessarily has inessential boundary in $S$. If $\Delta$ is the union of mutually disjoint s-discs for $S$, we can compress $S$ to a (possibly disconnected) surface $S|_{\Delta}$ by letting $S|_{\Delta}$ be the result of removing a regular neighborhood $N(\Delta)\approx(\Delta\times I)$ of $\partial\Delta$ from $S$ and then taking the union with $\Delta\times\partial I$. Observe that $S|_{\Delta}$ inherits a normal orientation from $S$ and that each disc in $\Delta$ leaves two scars on $S|_{\Delta}$.
@@
  • compressible if there is a compressing disc for $S$ and incompressible if there is no such disc;
  • weakly reducible if there exist disjoint s-discs $D_{1}$ and $D_{2}$ for $S$, with boundaries on the same component of $S$, such that near $S$, the discs $D_{1}$ and $D_{2}$ lie on opposite sides of the component of $S$ containing their boundaries and at least one of $D_{1}$ or $D_{2}$ is a compressing disc. (“Near SS” means that in a regular neighborhood $N(S)$ of $S$, which may be nonseparating in $M$, each $D_{i}\cap N(S)$ is a collar neighborhood of $\partial D_{i}$, and these collar neighborhoods lie on opposite sides of $S$);
@@
    There is a (possibly empty) collection $\Delta$ of mutually disjoint s-discs for $\partial_{+}C$, such that $C|_{\Delta}$ is the union of trivial compressionbodies and the designations of $\partial_{+}$ and $\partial_{-}$ are inherited from $C$. Such a collection $\Delta$ is a complete collection of s-discs for $C$.
A punctured trivial compressionbody is the result of removing the open regular neighborhood of a finite (possibly empty) collection of points from the interior of a trivial compressionbody. The positive boundary of a punctured product compressionbody is the same as the positive boundary prior to removing the open balls. Observe that a punctured product compressionbody is a compressionbody.
A vertical arc $\alpha$ in a compressionbody $C$ is a properly embedded arc such that there exists a complete collection $\Delta$ of s-discs for $C$ disjoint from $\alpha$ such that $\alpha$ is isotopic to a fiber in one of the product compressionbody components of $C|_{\Delta}$.
@@
    $C$ does not contain an s-disc for $\partial_{+}C$ if and only if it is a trivial compressionbody. If $C$ contains an s-disc for $\partial_{+}C$ but does not contain a compressing disc for $\partial_{+}C$, then it is a punctured product compressionbody.
@@
By definition of semi-compressing disc, $\partial D$ bounds at least one disc in $\partial_{-}C$. As every sphere in a compressionbody is separating and since $\partial_{-}C$ is incompressible in $C$, there is such a disc $E$ such that the sphere $D\cup E$ bounds a submanifold $W\subset C$ disjoint from $\partial_{-}C$. The result follows as in the proof of Lemma 3.1 (5). ∎
@@
Given any compressionbody $C$ with $\partial_{-}C\neq\varnothing$, there exists a collection of s-discs $D\subset C$ such that each component of $C|_{D}$ is either a handlebody or a product compressionbody and each component contains exactly one scar from $D$. Furthermore, if $A\subset C$ is the union of finitely many pairwise disjoint vertical annuli and compressing discs for $\partial_{+}C$, we may choose $D$ to be disjoint from $A$.
@@
[…]
Figure 3. By sliding 1-handles incident to $\partial_{+}C$, we may find discs separating a compressionbody into product compressionbodies and handlebodies so that each product compressionbody contains a single scar.
@@
where the unions appearing in the last two items are over all the components of $M\setminus\mathcal{H}$ and the equality holds after gluing. A generalized Heegaard surface $\mathcal{H}$ is oriented if each component of $\mathcal{H}\cup\partial M$ is assigned a transverse orientation such that for each compressionbody $\subset M\setminus\mathcal{H}, CC$ is a cobordism from $\partial_{-}C$ to $\partial_{+}C$ or vice-versa. We say that a compressionbody $C$ is above (resp. below) $\partial_{+}C$ if the transverse orientation for $\partial_{+}C$ points into (resp. out of) $C$. Two oriented generalized Heegaard surfaces are equivalent if they are ambient isotopic in $M$, via an orientation-preserving isotopy. We will only work with oriented generalized Heegaard surfaces in what follows.
@@
A generalized Heegaard surface is a Heegaard surface if and only if it is connected. In which case, $\mathcal{H}=\mathcal{H}^{+}$ and $\mathcal{H}^{-}=\varnothing$.